OpenStudy (anonymous):

Use the distributive property to simplify: −6(2u − 1)

OpenStudy (anonymous):

\[\huge -6 \times 2u + (-6) \times (-1) = -12u + 6\]

OpenStudy (anonymous):

what?

OpenStudy (diyadiya):

Remember this: \(\large a(b+c) = a*b + b*c\) did you understand this?

OpenStudy (anonymous):

kinda

OpenStudy (diyadiya):

Ok so same way −6(2u − 1) = ?

OpenStudy (anonymous):

\[\huge -a(b-c) = -ab + ac\]

OpenStudy (anonymous):

oh so -12-6?

OpenStudy (diyadiya):

Where is the u?

OpenStudy (anonymous):

ooh -12u-6...

OpenStudy (diyadiya):

Nope -6(2u-1) = -6*2u -1*-6 = 12u + 6 Its +6 because : when negative is multiplied by negative we get a positive number For example: -3 * -2 =6 -4 * -5 = 20 is this clear?

OpenStudy (anonymous):

yep thanks

OpenStudy (diyadiya):

Welcome :)